"Rocket Man" has a propulsion unit strapped to his back. He starts from rest on the ground, fires the unit, and accelerates straight upward. At a height of 18 m, his speed is 4.6 m/s. His mass, including the propulsion unit, has the approximately constant value of 128 kg. Find the work done by the force generated by the propulsion unit. (Neglect air resistance.)
WP =
Answer #1
given
intial speed = u = 0
final speed = v = 4.6m/s
final heigth = h = 18m
mass of the rocket man = m = 128kg

next_step.gif
intial energy = PE ( at ground) + 1/2mu2
                    = mg(0) + 1/2 * m * 0
                   = 0J
final energy = 1/2mv2 + mgh
                  = 2.39 * 104J
work done = final energy - intial energy
                   = 2.39 * 104J